# Treewidth of generalized Hamming graph, bipartite Kneser graph and generalized Petersen graph

### v1

@@ Acknowledgement
Case 4.3.1: If $k-s<0$, then $(k-1)-(s-1)<0$.
@@
Case 4.3.2: If $0\leq k-s\leq\left\lfloor((n-1)-t)/2\right\rfloor$, then $0\leq k-s\leq\left\lfloor(n-t-1)/2\right\rfloor\leq\left\lfloor(n-t)/2\right\rfloor$ and
@@
$A^{(2)}_{t,n-1,k,s}-A^{(2)}_{t-1,n-1,k-1,s-1}=-(\binom{n-s-1}{k+t-2s-1}-\binom{n-s-1}{k-s-1}),$
@@
$\leq 0.$
Therefore,
@@
$=$ $\binom{n-s-1}{k-s-1}-\binom{n-1}{k-1}\leq 0.$
Case 4.3.4: If $k-s>(n-1)-t$, then $k-s\geq n-t$.
@@
Step 5: verify the base case of induction.
@@
$\sum_{q=1}^{t-1}\binom{n}{k+q}+r(M^{(t,n)}_{k,k+t})$ $=\sum_{q=0}^{t-1}\binom{n}{k+q}+\sum_{a=0}^{\left\lfloor(n-t-1)/2\right\rfloor}\left(\binom{t+2a}{t+a-1}-\binom{t+2a}{a-1}.\right)$ (21)
@@
From definition, when $k=0$, we have that $\widetilde{r}(M^{(t,n)}_{k,k+p})$ reach its maximal when $p=t$. Similarly, when $k+p=n$, $\widetilde{r}(M^{(t,n)}_{k,k+p})$ reach its maximal when $k=n-t$, that is, $p=t$. When $k>0$ and $k+p<n$, we have $\widetilde{r}(M^{(t,n)}_{k,k+p})\leq\widetilde{r}(M^{(t,n)}_{k-1,k+p+1})$ when $p+2\leq t$. Hence, we only need to prove $\widetilde{r}(M^{(t,n)}_{k,k+p})$ is no larger than Eq 21 for the following two cases:
@@
Case 1: $p=t-1$. In this case, we have $\widetilde{r}(M^{(t,n)}_{k,k+p})=\widetilde{r}(M^{(t-1,n)}_{k,k+p})\leq bw(M^{(t-1,n)})$ by definition. Since the value of $bw(M^{(t-1,n)})$ can be calculated from Lemma 3.11 by induction, and it is not hard to verify the RHS of Eq. 14 is increasing with respect to $t$. Then we can reach our conclusion in this case.

### v3

@@ Acknowledgement
First, if t=n2t=n-2, then
$r^{(4)}_{t,n,k,t-2s}$ $=\binom{n-1}{k+t-2s-1}+r(M^{(t,n-1)}_{k,k+t-2s})$
$=\binom{n-1}{k+t-2s-1}+\binom{n-1}{k}+\binom{n-1}{k+t-2s}-1.$
Moreover, we have
$r^{(2)}_{t,n,k,t-2s}$ $=\binom{n-1}{k}+\binom{n-1}{k+t-2s-1}+r(M^{(t-1,n-1)}_{k-1,k+t-2s})$
If ks<=0k-s<=0 or ks>=nt=2k-s>=n-t=2, then $r(M^{(t-1,n-1)}_{k-1,k+t-2s})\geq D_{t-1,n-1,k-1,s-1}=\binom{n-1}{k+t-2s}-1$. Otherwise, ks=1k-s=1 and then $r(M^{(t-1,n-1)}_{k-1,k+t-2s})\geq\binom{n-1}{k-1}=\binom{n-1}{k+t-2s-1}$. In both cases, we can derive $r^{(4)}_{t,n,k,t-2s}\leq r^{(2)}_{t,n,k,t-2s}$. Thus, in the following, we assume t<n2t<n-2.
Case 4.3.1: If ks<0k-s<0, then (k1)(s1)<0(k-1)-(s-1)<0.
@@
Case 4.3.2: If $0\leq k-s\leq\left\lfloor((n-1)-t)/2\right\rfloor$, then $0\leq k-s\leq\left\lfloor(n-t-1)/2\right\rfloor\leq\left\lfloor(n-t)/2\right\rfloor$ and $2(k-s)\leq n-t-1$. We can derive that
@@
$A^{(2)}_{t,n-1,k,s}-A^{(2)}_{t-1,n-1,k-1,s-1}=-\left(\binom{n-s-1}{k+t-2s-1}-\binom{n-s-1}{k-s-1}\right)$,
@@
$=-\binom{2n-2k-t+s-2}{n-k-2}+\binom{2n-2k-t+s-2}{n-t-k+s-2}$.
Write $\lambda\triangleq k-s-\frac{n-t}{2}$, then $\lambda\geq 0$ by the condition. Therefore,
@@
$\leq$ $\binom{n-s-1}{k+t-2s-1}-\binom{n-1}{k-1}-\binom{2n-2k-t+s-2}{n-k-2}+\binom{2n-2k-t+s-2}{n-t-k+s-2}$
$\leq$ $\binom{n-s-1}{k-s-2\lambda}-\binom{n-1}{k-1}-\binom{n-s-2-2\lambda}{k-s-2\lambda}+\binom{n-s-2-2\lambda}{k-s-2\lambda-2}$.
If $\lambda=0$, it is easy to verify $r^{(4)}_{t,n,k,t-2s}-r^{(2)}_{t,n,k,t-2s}\leq 0$ with the above equation. If $\lambda\geq 1$, then we first claim that $\binom{n-s-1}{k-s-2\lambda}\leq\binom{n-s-1}{k-s-1}$. If $k-s-1\leq\frac{n-s-1}{2}$, it holds naturally. Otherwise, when $k-s-1>\frac{n-s-1}{2}$, note that $k-s-2\lambda\leq\frac{n-s-1}{2}$ holds, since $t\geq 2s\geq 2$. It suffices to prove that $k-s-1-\frac{n-s-1}{2}\leq\frac{n-s-1}{2}-(k-s-2\lambda)$ which is easy to verify. As a consequence, we have
$r^{(4)}_{t,n,k,t-2s}-r^{(2)}_{t,n,k,t-2s}$
$\leq$ $\binom{n-s-1}{k+t-2s-1}-\binom{n-1}{k-1}$
$=$ $\binom{n-s-1}{k-s-2\lambda}-\binom{n-1}{k-1}$
$\leq$ $\binom{n-s-1}{k-s-1}-\binom{n-1}{k-1}\leq 0.$
Case 4.3.4: If ks>(n1)tk-s>(n-1)-t, then $k-s\geq n-t$.
@@
Step 5: verify Eq 13 when $k>0$, k+t2s<nk+t-2s<n and $s=0$. In this case, $r^{(2)}_{t,n,k,t-2s}$$\infty$ and ${\{r^{(i)}_{t,n,k,t-2s}\}}_{1=1,3,4}$ are positive.
Case 5.1: If $1\leq k<\lfloor(n-t)/2\rfloor$, then we have $k\leq\lfloor((n-1)-t)/2\rfloor$ and $2k\leq n-t-1$. We first verify that $r^{(4)}_{t,n,k,t-2s}$ is always equal to RHS of Eq 13.
$r^{(4)}_{t,n,k,t-2s}$ $=\binom{n-1}{k+t-2s-1}+r(M^{(t,n-1)}_{k,k+t-2s})$
$=\binom{n-1}{k+t-2s-1}+\binom{n-1}{k}+A^{(1)}_{t,n-1,k,s}+A^{(2)}_{t,n-1,k,s}+C_{t,n-1,k,s}.$
Note that $A^{(1)}_{t,n-1,k,s}=A^{(1)}_{t,n,k,s}$, $C^{(1)}_{t,n-1,k,s}=0=C_{t,n,k,s}$ when $s=0$, and
$A^{(2)}_{t,n-1,k,s}$ $=A^{(2)}_{t,n,k,s}-\left(\binom{n-1}{k+t-2s-1}-\binom{n-1}{k-s-1}\right)$.
Then we can derive that
$r^{(4)}_{t,n,k,t-2s}$ $=\binom{n}{k}+A^{(1)}_{t,n,k,s}+A^{(2)}_{t,n,k,s}+C_{t,n,k,s}.$
Consequently, $r^{(4)}_{t,n,k,t-2s}$ is always equal to RHS of Eq 13 and we then only need to prove $r^{(1)}_{t,n,k,t-2s}\leq r^{(4)}_{t,n,k,t-2s}$ and $r^{(3)}_{t,n,k,t-2s}\leq r^{(4)}_{t,n,k,t-2s}$. We omit the details here.
Case 5.2: If $\lfloor(n-t)/2\rfloor<k\leq n-t$, then $\lfloor(n-1-t)/2\rfloor\leq k-1\leq(n-1)-t$. We first verify that $r^{(1)}_{t,n,k,t-2s}$ is always equal to RHS of Eq 13.
$r^{(1)}_{t,n,k,t-2s}$ $=\binom{n-1}{k}+r(M^{(t,n-1)}_{k-1,k+t-2s-1})$
$=\binom{n-1}{k}+\binom{n-1}{k-1}+B_{t,n-1,k-1,s}+C_{t,n-1,k-1,s}$
$=\binom{n}{k}+B_{t,n,k,s}+C_{t,n,k,s}.$
Consequently, $r^{(1)}_{t,n,k,t-2s}$ is always equal to RHS of Eq 13 and we then only need to prove $r^{(3)}_{t,n,k,t-2s}\leq r^{(1)}_{t,n,k,t-2s}$ and $r^{(4)}_{t,n,k,t-2s}\leq r^{(1)}_{t,n,k,t-2s}$. We omit the details here.
Step 6: verify the base case of induction.
@@
[…]
@@
From definition, when $k=0$, we have that $\widetilde{r}(M^{(t,n)}_{k,k+p})$ reach its maximal when $p=t$. Similarly, when $k+p=n$, $\widetilde{r}(M^{(t,n)}_{k,k+p})$ reach its maximal when k=ntk=n-t, that is, $p=t$. When $k>0$ and $k+p<n$, we have $\widetilde{r}(M^{(t,n)}_{k,k+p})\leq\widetilde{r}(M^{(t,n)}_{k-1,k+p+1})$ when $p+2\leq t$. Hence, we only need to prove $\widetilde{r}(M^{(t,n)}_{k,k+p})$ is no larger than Eq 21 for the following two cases:
@@
Case 1: p=t1p=t-1. In this case, we have $\widetilde{r}(M^{(t,n)}_{k,k+p})=\widetilde{r}(M^{(t-1,n)}_{k,k+p})\leq bw(M^{(t-1,n)})$ by definition. Since the value of $bw(M^{(t-1,n)})$ can be calculated from Lemma 3.12 by induction, and it is not hard to verify the RHS of Eq. 14 is increasing with respect to $t$. Then we can reach our conclusion in this case.